OpenStudy (anonymous):

is this simplified? [-26+ - 2sqrt(163)]/-6 or could I simplify it to [-13+ - sqrt(163)]/-3 or does the negative cancel out so it would be [13 + - sqrt(163)]/3??

OpenStudy (anonymous):

i hate simplifying. I never know when I am done!

OpenStudy (ash2326):

We have \[\frac{-26+-2\sqrt{163}}{-6}\] first cancel 2 from numerator and denominator \[\frac{-13+-\sqrt{163}}{-3}\]now this can be written as \[\frac{-13-\sqrt{163}}{-3}\] remove" - "from all the terms \[\frac{13+\sqrt{163}}{3}\]

OpenStudy (anonymous):

ok. thank you, but wouldn't the "+-" still be there? not just "+"

OpenStudy (ash2326):

\[+ \times -=-\]

OpenStudy (anonymous):

but its still a square root, which could be both positive or negative numbers

OpenStudy (ash2326):

oops I got your point then \[\frac{13\pm\sqrt{163}}{3}\]

OpenStudy (ash2326):

yeah you're right :) @a.galvan2

OpenStudy (anonymous):

:)

OpenStudy (campbell_st):

aren't there still 2 answers... as it can be split \[\frac{-26}{-6} \pm \frac{\sqrt{163}}{-6} = \frac{13}{3} \mp \frac{\sqrt{163}}{3}\]